What is the (a) H+ ion concentration, (b) pH, (c) OH- ion concentration and (d) pOH of a 0.001 M solution of NaOH?


1
Expert's answer
2022-03-16T04:54:04-0400

NaOH is a strong base because NaOH completely dissociates into sodium (Na+) and hydroxide (OH) ions in aqueous solutions. 

Equation representing the dissociation of NaOH:

NaOH(aq) → Na+(aq) + OH(aq)


The molarity of NaOH is 0.001 M

CM(NaOH) = [OH] = [Na+] = 0.001 M

Thus,

[OH] = 0.001 M

pOH = −log[OH] = −log[0.001] = 3

pH = 14 − pOH = 14 − 3 = 11

[H+] = 10−pH = 10−11 = 1.0×10−11 M


Answer:

(a): [H+] = 1.0×10−11 M

(b): pH = 11

(c): [OH] = 0.001 M

(d): pOH = 3